OpenStudy (anonymous):

How do you differentiate y=π√π(x^4 +2x)e^2x

OpenStudy (anonymous):

ignore the \[\pi\sqrt{\pi}\] and take the derivative of \[(x^4+2x)e^{2x}\] using the product rule \[(fg)'=f'g+g'f\] with \[f(x)=x^4+2x,f'(x)=4x^3+2, g(x)=e^{2x}, g'(x)=2e^{2x}\] plug into formula for the product rule, multiply by the constant at the end